OpenStudy (anonymous):

What is the value of g? -25/g=25 g =

OpenStudy (solomonzelman):

multiply both sides g, divide both sides by 25.

OpenStudy (solomonzelman):

times* g

OpenStudy (anonymous):

Okay.....

OpenStudy (anonymous):

A lil confusing, dont we have to cancel out both sides?

OpenStudy (solomonzelman):

yes, you will cancel out g on the left side when you multiply both sides times g. Then, when you divide both sides by 25 you will cancel 25 on the right side.

OpenStudy (anonymous):

Wont that be 0?

OpenStudy (solomonzelman):

If you divide and multiply by non-zero numbers you should be getting a zero.

OpenStudy (solomonzelman):

-25/g = 25 multiply both sides times g, (-25/g) • g = 25 • g the g on the left side cancels, and you get: -25 = 25 • g multiply both sides times 25,

OpenStudy (solomonzelman):

from there, proceed on your own please.

OpenStudy (anonymous):

Okay thank you

OpenStudy (anonymous):

i got 625

OpenStudy (solomonzelman):

I made an error sorry

OpenStudy (solomonzelman):

you divide by 25, so that you cancel 25 on the right, to isolate the g.

OpenStudy (solomonzelman):

-25/g = 25 multiply both sides times g, (-25/g) • g = 25 • g the g on the left side cancels, and you get: -25 = 25 • g divide by 25, (-25) / 25 = 25g / 25 -1 = g

OpenStudy (anonymous):

Okay thank you :)

OpenStudy (solomonzelman):

Yw